1. Calculate labor force participation rate in an economy with an adult population of 100, 50 of whom hold jobs, 10 of whom are looking for work, and 15 of whom are retired.

2. Calculate a) labor force, b) labor force participation rate, and c) unemployment rate if the population of a country is 380 million people out which 92 million are under the age of 16, 58 million don't want to work and 20 million are looking for work.

3. Calculate the inflation rate in 2005 if the index number in 2005 was 120 and the index number in 2004 was 135.
1
Expert's answer
2018-03-29T11:39:07-0400
1)labor force participation rate=(labor force/adult population)*100%=((50+10)/100)*100%=60%
2)Labor force=380-92-58=230million
participation rate=(230/(380-92))*100%=79.86
unemployment rate=20/230*100%=8.695%
3) inflation rate =(135-120)/120*100=12,5%